OpenStudy (anonymous):

Please help! :) A spacecraft which weighed 1,500.0 pounds on earth is traveling in deep space at a constant velocity of 100,000.0 ft/sec. Ground control wants to initiate a rocket burn that will accelerate the spacecraft 500.0 ft/sec^2. What is the mass in slugs of the spacecraft? What is the force of that rocket burn?

OpenStudy (anonymous):

\[1 slug=1\frac{ pound \times second ^{2} }{ ft }\] Thus if you know the acceleration due to gravity on Earth (g) you can solve for the mass:\[m=\frac{ 1500 }{ g }\] Then, knowing the mass, you can calculate the force required. We know that \[\frac{ d }{ dt }v(t)=0\] so there is no acceleration. Remember that \[\Delta F= m \times \Delta a\] if there is no change in mass. Thus since you know that the change in acceleration is 500, just solve for the force. Don't forget your units!

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're welcome.